OpenStudy (anonymous):

How do I set up the polar or trig form of -81i ?

OpenStudy (anonymous):

the angle is \(270\) or \(\frac{3\pi}{2}\) as you are on the \Ii\) axis

OpenStudy (anonymous):

so it is \(81(\cos(\frac{3\pi}{2}+i\sin(\frac{3\pi}{2}))\)

OpenStudy (anonymous):

thank you!